OpenStudy (anonymous):

using substitution, what is the system of the equation -4y-3x=-11 x-2y=17

OpenStudy (anonymous):

-4y - 3x = -11 and x - 2y =17 from 2nd equation, x = 17 + 2y, substitute this into the 1st equation -4y - 3(17+2y) = 17 now you can solve for y

OpenStudy (anonymous):

thankyou